40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 18:26:05 2019, gautam, Update, IOO, IPPOS recommissioned IMG_7330.JPGScreenshot_from_2019-02-20_19-43-27.png
    Reply  Fri Feb 22 12:19:46 2019, gautam, Update, IOO, TT coil driver Vmon 
Message ID: 14469     Entry time: Fri Feb 22 12:19:46 2019     In reply to: 14467
Author: gautam 
Type: Update 
Category: IOO 
Subject: TT coil driver Vmon 

To debug the issue of the suspected drifting TTs further, I temporarily hijacked CH0-CH8 of ADC1 in the c1lsc expansion chassis, and connected the "MON" outputs of the coil drivers (D010001) to them via some DB9 breakouts. The idea is to see if the problem is electrical. We should see some  slow drift in the voltage to the TTs correlated with the spot walking off the IPPOS QPD. From the wiring diagram, it doesn't look like there is any monitoring (slow or fast) of the control voltages to the TT coils, this should be factored into the Acromag upgrade of c1iscaux/c1iscaux2. EPICS monitoring should be sufficient for this purpose so I didn't setup any new DQ channels, I'll just look at the EPICS from the IOP model.

Already, in the last ~1 hour, there has been considerable drift - see Attachment #2. The spot, which started at the center of the CCD monitor, has now nearly drifted off the top end. The ITMX and BS Oplev spots have been pretty constant over the same timescale, so it has to be the TTs?
ELOG V3.1.3-